OpenStudy (anonymous):

Liz has 2 red, 2 white, and 3 blue marbles in a cup. She draws two marbles at random and does not replace the first one. What is the probability that she will draw a white marble and then a blue marble or a blue then a red?

OpenStudy (sasogeek):

ok so umm there are a total of 7 marbles. 2 red 2 white 3 blue the probability that she draws a white marble is 2/7 , if she picks a white marble but doesn't replace it, the total number of marbles left are 6. so the probability that she would then pick a blue marble will be 3/6. now she's not picking white OR blue, she's picking white AND blue... so our equation will be \(\large P(W_{1}nB_{2}) =\frac{2}{7}\times \frac{3}{6} = \frac{2}{14} \)

OpenStudy (anonymous):

the key is 2/7

OpenStudy (sasogeek):

well i think i didn't answer the whole question lol, i forgot the last part

OpenStudy (anonymous):

I think you not finish

OpenStudy (sasogeek):

if she picks a blue, then a red, the probability will be \(\large P(B_{1}nR_{2})=\frac{3}{7}\times \frac{2}{6}=\frac{1}{7} \) not that she's picking white then blue OR blue then red this means that we have to ADD the 2 probabilities so that would be \(\large \frac{1}{7}+\frac{1}{7}=\frac{2}{7} \) note that \(\large \frac{2}{14}=\frac{1}{7} \)

OpenStudy (anonymous):

ok,ty

OpenStudy (sasogeek):

you're welcome :) glad i could help :)

OpenStudy (sasogeek):

btw i meant to say that "NOTE that she's picking white then ..." and not "NOT that she's picking white then ..." :)